OpenStudy (anonymous):

Decrypt the following cipher text obtained from auto key cipher by using exhaustive key search MALVVMAFBHBUQPTSOXALTGVWWRG

OpenStudy (anonymous):

I tried different keys but I got exhausted any help will be appreciated

OpenStudy (anonymous):

So the way is Convert the alphabetic string to the numeric string M A L V V M A F B H B U Q P T S O X A L T G V W W R G 12 1 11 21 21 12 1 5 2 7 2 20 16 15 19 18 14 23 1 11 19 6 21 22 22 17 6 12 1 11 21 21 12 1 5 2 7 2 20 16 15 19 18 14 23 1 11 19 6 21 22 22 17 6 Then we compute X1 = d9 (12) = (12 - 8 ) mod 26 = 3 X2 = d3(1) = 24 X3 = d24(11) = 13 x4 = d14(21) = 7 x5 = d7(21) = 14 x6 = d14(12) = 24

OpenStudy (anonymous):

You can use this calculator if you are lazy http://britton.disted.camosun.bc.ca/modarith/modular/modcalc.html

OpenStudy (anonymous):

The above trial we assume that the key is 9 we start with

OpenStudy (anonymous):

Need help in finding the word and the key used ??

OpenStudy (anonymous):

If we assume that the key k = 8 then X1 = d8 (12) = (12 - 8 ) mod 26 = 4 X2 = d4(1) = 23 X3 = d23(11) = 14 x4 = d14(21) = 7 x5 = d7(21) = 14 x6 = d14(12) = 24 x7 = d24(1) = 3 EXOHO

OpenStudy (anonymous):

Not sure if I did it right or I overlooked some thing

OpenStudy (bahrom7893):

i dont understand this one.. what do u mean by all those numbers?

OpenStudy (anonymous):

Ok you first convert the letters above to numbers you know numeric string

OpenStudy (bahrom7893):

nope, never heard of the numeric string.

OpenStudy (anonymous):

lool you know it bahrom :P then you compute for the first letter lets say assuming Key = 9 its X1 = d9(M) = d9 (12) = (12-9) mod 26 = 3 you can use the calculator on the link above

OpenStudy (anonymous):

then take that result 3 we sub to get the second letter X2 = d3(A) = d3(1) = 24

OpenStudy (anonymous):

after finding all the X numbers then we convert them back to the letters that corresponds to those numbers example X1 = 3 which is C

OpenStudy (anonymous):

X2 = 24 which is Y

OpenStudy (anonymous):

and so on then if that doesnt work we use a different key K = 8 or K = 10 and run the same operation

OpenStudy (bahrom7893):

im totally lost

OpenStudy (anonymous):

come on bahrom

OpenStudy (anonymous):

re read it again cant figure out the key

OpenStudy (anonymous):

MALVVMAFBHBUQPTSOXALTGVWWRG

OpenStudy (bahrom7893):

gimme like 15 mins.

OpenStudy (bahrom7893):

i was helping someone before we started, let me jstu finish that.

OpenStudy (anonymous):

Sure

OpenStudy (bahrom7893):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when all else fails, brute force attack...

OpenStudy (anonymous):

It works when K = 19 do you believe that